OpenStudy (anonymous):

What is the equation of the line, in slope-intercept form, that passes through (-1, 3) and (2, -3)? y= -2x+ 1 y= -3x+ 7 2x+y- 1 = 0

OpenStudy (anonymous):

slope=-3-3/2+1=-6/3=-2

OpenStudy (anonymous):

A